Meeting on

”QUATERNIONIC STRUCTURES IN MATHEMATICS AND PHYSICS” Trieste, 5-9 September 1994

FINAL PROGRAMME

Monday, 5 September 1994

11:00  Opening
11:15-12:00 A. LICHNEROWICZ  Espaces Homogènes de contact

complexes et structures qua- -ternioniques

Lunch Break

15:30-16:15 L. RYAN Conformal structure in Clif-

ford analysis

17:00-17:25 T. NITTA  Vanishing theorems over qua-

-ternion-Kähler manifolds

17:30-17:55 Y. NAGATOMO  Instantons on quaternion-Käh-

-ler manifolds

Tuesday, 6 September 1994

9:30-10:15 B. DE WIT Special geometry and quater-

nionic manifolds

11:00-11:45 L. BERARD-BERGERY Some remarks on Riemannian

homogeneous spaces and fibrations

12:00-12:45 E. BONAN  More on quaternions

Lunch Break

15:30-15:55 V. CORTES  Alekseevskian spaces

16:00-16:25 R. GOTO On hyper-Kähler manifolds of type A_{∞}, D_{∞} and elliptic fibrations

17:00-17:25 T. HANGAN  Transformations infinitésima-

-les homographiques des espa- -ces Riemanniens symétriques compacts Wednesday, 7 September 1994

9:30-10:15 A. VAN PROEYEN Isometries of special Kähler and

quaternionic manifolds

11:00-11:45 A. SWANN  Quaternionic Kähler metrics and

nilpotent orbits for classical groups

12:00-12:25 S. LUGOJAN  About quaternionic derivative
12:30-13:00 P. PICCINNI  Locally conformal quaternion

Kähler structures

Lunch Break

Afternoon free

Thursday, 8 September 1994

9:30-10:15 A. FUJIKI  Nagata threefold and twistor spa-

-ces

11:00-11:45 O. HIJAZI  Twistor operators on spin quater-

-nionic manifolds

12:00-12:45 L. PERNAS  On operators linked with quater-

-nionic structures. Applications to the hyperbolic quaternionic spaces.

Lunch Break

15:00-15:25 J. CNOPS  Dirac operators and Stokes theo-

-rem

15:30-15:55 U. SEMMELMANN Complex contact structures and the first eigenvalue of the Dirac operator on Kähler manifolds

16:00-16:25 R. I. TORRES On the canonical homology of a

compact symplectic manifold

17:00-17:25 L. UGARTE A corresponding Dolbeaut cohomology for G_{2} manifolds Friday, 9 September 1994

9:30-10:15 K. GALICKI  On 3-Sasakian manifolds

11:00-11:45 V. OGIEVETSKY Self-dual gravity and supergravities. Supersymmetrization of quaternionic structures

12:00-12:45 T. NAGANO  Symmetric spaces and quater-

-nionic structures